### v1

@@ 3.2 Proof of main theorem
We proceed with the proof following the aforementioned four steps. First, we need to demonstrate how the convergence target $\Lambda_{t}$ can be expressed in terms of the operator $\mathcal{G}^{\alpha-1}$ in (3.2).
@@
The second step, we prove that $Y^{(n)}\Rightarrow Y$ in Hölder space $\mathcal{C}^{\lambda}$ for $\lambda<1/2$. The Donsker invariance principle for convergence in Hölder spaces was first established by Lamperti, (1962). Račkauskas and Suquet, (2004) provided necessary and sufficient conditions for the convergence of i.i.d. partial sum processes in Hölder spaces, while results for martingale difference sequences were derived by Giraudo, (2016) from a dynamical systems perspective.
@@
Note that
$Y_{t}^{(n)}=\theta t+\frac{\theta}{n\omega_{n}}\left(\sum_{i=0}^{\lfloor nt\rfloor-1}\Delta M^{(n)}_{i}+(nt-\lfloor nt\rfloor)\Delta M^{(n)}_{\lfloor nt\rfloor}\right)$,
by Theorem 2.2 in Giraudo, (2016), it suffices to prove that
$\lim_{t\to\infty}t^{p}\mathbb{P}\left(|\Delta M_{k}^{(n)}|>t\right)=0\ \text{and}\ \mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2}|\mathcal{F}_{k-1}\right]\in L^{p/2}.$ (3.8)
Then
$n^{-1/2}\left(\sum_{i=0}^{\lfloor nt\rfloor-1}\Delta M^{(n)}_{i}+(nt-\lfloor nt\rfloor)\Delta M^{(n)}_{\lfloor nt\rfloor}\right)\Rightarrow\sigma W_{t}\ \text{in}\ \mathcal{C}^{1/2-1/p}([0,1]),$
where $\sigma^{2}=\lim_{n\to\infty}\frac{1}{n}\mathbb{E}[\sum_{i=0}^{n-1}\Delta M^{(n)}_{i}]^{2}\in L^{1}$. For $p^{\prime}>0$, if $\mathbb{E}|\Delta M_{k}^{(n)}|^{p^{\prime}}<\infty$, by Markov’s inequality, for any positive $p<p^{\prime}$, we have
$t^{p}\mathbb{P}\left(|\Delta M_{k}^{(n)}|>t\right)\leq t^{p-p^{\prime}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p^{\prime}}\to 0,\text{as}\ t\to\infty.$
By Jensen’s inequality,
$\mathbb{E}\left(\mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2}|\mathcal{F}_{k-1}\right]\right)^{\frac{p}{2}}\leq\mathbb{E}\left(\mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2\cdot\frac{p}{2}}|\mathcal{F}_{k-1}\right]\right)=\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
Therefore, Assumption 3 ensures that condition (3.8) holds for any $p>2$. Moreover,
$\lim_{n\to\infty}\frac{1}{n}\mathbb{E}\left[\sum_{i=0}^{n-1}\Delta M^{(n)}_{i}\right]^{2}=\lim_{n\to\infty}\sum_{i=0}^{n-1}\frac{1}{n}\mathbb{E}\left[\mathbb{E}\left(|\Delta M_{i}^{(n)}|^{2}|\mathcal{F}_{i-1}\right)\right]=\lim_{n\to\infty}\sum_{i=0}^{n-1}\frac{1}{n}\mathbb{E}\left[U^{2}(\lambda_{i}^{(n)})\right]=\gamma^{2},$
together with $\frac{\theta}{\sqrt{n}\omega}\to\nu$, we have the weak convergence
$Y_{t}^{(n)}\Rightarrow\theta t+\nu\gamma W_{t}=Y_{t}$
in $\mathcal{C}^{1/2-1/p}([0,1])$, for any $p>2$. ∎
Subsequently, we establish the continuity of the operator $\mathcal{G}^{\alpha-1}$, thereby enabling the application of the continuous mapping theorem. Indeed, this result is encompassed within Proposition 2.2 of Horvath et al., (2024), where a proof of continuity for general GFOs is provided. We directly apply their result to obtain the following proposition.
@@
According to Lemma 3.10 in Horvath et al., (2024), we can readily extend this weak convergence to the continuous function space $\mathcal{C}([0,1])$ and Skorokhod space $\mathcal{D}([0,1])$.
We have established the convergence of $\widetilde{\Lambda}_{t}^{(n)}$. However, our primary process of interest is $\Lambda_{t}^{(n)}$ given by (2.5), with $\widetilde{\Lambda}_{t}^{(n)}$ serving merely as an auxiliary process. Finally, we will demonstrate that the two processes are asymptotically indistinguishable. By virtue of Theorem 3.1 in Billingsley, (2013), it follows that $\Lambda_{t}^{(n)}\Rightarrow\Lambda_{t}$.
@@
By comparing (2.5) with (3.4), we have
It is evident that $I_{1}\to 0$. By (3.6) and Dini’s theorem, $I_{2}$ converges to zero uniformly. As for $I_{3}$,
$\mathbb{E}|\sup_{t\in[0,1]}I_{3}|\leq\frac{\theta}{n\omega_{n}}\mathbb{E}\left|\Delta M_{1}^{(n)}\right|\sup_{s\in(0,t-t_{\lfloor nt\rfloor}]}\left|f^{\alpha,\kappa}(s)\right|\leq\frac{C}{n\omega_{n}}\left(\frac{1}{n}\right)^{\alpha-1}\sim n^{1/2-\alpha}\to 0.$
Finally, we focus on $I_{4}$, aiming to prove that its supremum converges to zero in probability. Noting that $n(1-a_{n})\psi_{\lfloor nt\rfloor}^{(n)}/\rho_{n}(t)=a_{n}\to 1$, we will henceforth replace $n(1-a_{n})\psi_{\lfloor nt\rfloor-i}^{(n)}$ with $\rho_{n}(t-t_{i})$. Then, we employ the left-endpoint rule to approximate the integral $\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s$, and this approximation is uniformly since $f^{\alpha,\kappa}\in L^{r}$ for $r<1/(1-\alpha)$.
Therefore, for any $\varepsilon>0$,
$\begin{split}\mathbb{P}\left(\sup_{t\in[0,1]}|I_{4}|\geq\varepsilon\right)&\sim\mathbb{P}\left(\sup_{k=1,\ldots,n}\left|\frac{\theta}{n\omega_{n}}\sum_{i=0}^{k-1}\left[\rho_{n}\left(t_{k}-t_{i}\right)-f^{\alpha,\kappa}(t_{k}-t_{i})\right]\Delta M_{i}^{(n)}\right|\geq\varepsilon\right)\\ &\leq\left(\frac{\theta}{\varepsilon n\omega_{n}}\right)^{2}\mathbb{E}\left|\sum_{i=0}^{n-1}\left[\rho_{n}(t_{k}-t_{i})-f^{\alpha,\kappa}(t_{k}-t_{i})\right]\Delta M_{i}^{(n)}\right|^{2}\\ &=\left(\frac{\theta}{\varepsilon n\omega_{n}}\right)^{2}\sum_{i=0}^{n-1}\left[\rho_{n}(t_{k}-t_{i})-f^{\alpha,\kappa}(t_{k}-t_{i})\right]^{2}\mathbb{E}\left[\Delta M_{i}^{(n)}\right]^{2}\\ &=\frac{\theta^{2}\sup_{n}\mathbb{E}\left[\Delta M_{i}^{(n)}\right]^{2}}{\varepsilon^{2}n\omega_{n}^{2}}\sum_{i=1}^{n}\left[\rho_{n}(t_{i})-f^{\alpha,\kappa}(t_{i})\right]^{2}\frac{1}{n}\\ &\lesssim C_{\varepsilon}\|\rho_{n}-f^{\alpha,\kappa}\|_{2}^{2}\end{split}$
By Proposition 3, it follows that the above probability converges to zero. Consequently, $\Lambda_{t}^{(n)}$ and $\widetilde{\Lambda}_{t}^{(n)}$ are asymptotically indistinguishable. ∎
@@ 4.1 Gamma-GED-EGARCH($\infty$) model
If $\varepsilon_{n}$ has unit variance, $\sqrt{\varphi(\lambda_{n})}$ is in fact the conditional volatility of $y_{n}$. Here, the function $\varphi:\mathbb{R}\to\mathbb{R}^{+}$ is monotonic and differentiable, referred to as the link function. Moreover, if $\varphi$ is the identity mapping (restricted to $\mathbb{R}^{+}$), then the time-varying parameter $\lambda_{n}$ directly characterizes the conditional variance of $y_{n}$. According to Buccheri et al., (2021), the score in this case can be expressed as
@@
Following the terminology in Harvey and Lange, (2017), we refer to this model as the Gamma-GED-EGARCH($\infty$). When $\nu=2$, the GED reduces to standard normal distribution, and this model corresponds to the EGARCH model in Nelson, (1991) but without asymmetry. Based on the previous definitions, we define the following sequence scaling process:
$\begin{split}&X_{t}^{(n)}=\frac{1}{\sqrt{n}}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}\varepsilon_{i},\quad\varepsilon_{i}|\mathcal{F}_{i-1}\stackrel{{\scriptstyle d}}{{\sim}}\ \text{GED}(\nu),\\ &\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\frac{\nu}{2}|\varepsilon_{n-i}|^{\nu}-1\right],\end{split}$ (4.1)
where $\phi_{i}\sim K/i^{1+\alpha}$, $1-a_{n}\sim n^{-\alpha}$, $\omega_{n}\sim n^{-1/2}$.
@@
Denote $\mathbb{E}_{k-1}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{k-1}]$. By the property of GED, for any $p\geq 2$,
Thus,
$\mathbb{E}[|\nabla_{k}|^{p}]=\sum_{i=0}^{p}\binom{k}{i}(-1)^{k-i}\left(\nu/2\right)^{i}\mathbb{E}[|\varepsilon_{k}|^{\nu i}]<\infty.$
Especially, it can be easily computed that
The second equation then follows from Theorem 1. From the fact that $\varepsilon$ with its score are naturally uncorrelated, by Donsker’ invariance principle,
$\left(\sum_{i=1}^{\lfloor nt\rfloor}\frac{\varepsilon_{i}}{\sqrt{n}},\sum_{i=1}^{\lfloor nt\rfloor}\frac{\nabla_{i}}{\sqrt{n}}\right)\Rightarrow(B_{t},W_{t})$
are two independent Brownian motions. ∎
@@ 4.2 A extension as rough volatility approximations
The limit (4.2) is actually a rough volatility model, where $\Lambda_{t}$ characterizes the log-volatility of assets log-price $X_{t}$. Note that $\Lambda_{t}$ follows a rough OU process but is different with that in Gatheral et al., (2018) which characterizes log-volatility with a OU process driven by rough Brownian motion. Because in this volatility process (4.2), the kernel $(t-s)^{\alpha-1}$ not only emerge in noise but also in drift term. This setting is analogous with rough CIR in Jaisson and Rosenbaum, (2016) or rough Heston in El Euch and Rosenbaum, (2019).
However, before it can be considered a “useful” rough volatility model, there are still two important issues we need to deal with. First, the initial value of $(X_{t},\Lambda_{t})$ is zero in (4.2); second, the two Brownian motions are independent, which prevents the model from capturing the leverage effect in the market.
For the first one, we adopt the strategy in Wang and Cui, (2025), add a baseline in the dynamic of $\lambda_{t}^{(n)}$, that is
@@ 4.3 Numerical simulation
Based on the convergence result, we will simulate the time series (4.3) to approximate rough volatility model (4.4). Accordingly, the simulated paths can be employed to price a variety of options via Monte Carlo methods. In this part, the parameter choices are aligned with those adopted in prior numerical studies, such as Callegaro et al., (2021), Ma and Wu, (2022) and Wang and Cui, (2025), these values are:
@@
It can be observed that as $\alpha$ increases, the volatility paths exhibit greater smoothness, and the mean-reverting behavior becomes more evident. In other words, a smaller $\alpha$ not only implies rougher paths but also corresponds to a slower mean-reversion, which is consistent with the observations reported in Gatheral et al., (2018).
@@
Table 1 reports the option prices computed via Algorithm 1, with standard errors shown in parentheses. We set $M=10^{5}$, $n=500$, $T=1$, with strike prices $K=80,90,100,110,120$. For the barrier options, the barrier level is set to $B_{u}=110$ for the Up-In call and $B_{d}=90$ for the Down-Out put.
Table 1: Option pricing results under the rough volatility model (4.4)
@@
To accelerate the path generation process, we employ two techniques: parallel computing and FFT-based convolution. Parallel computing distributes the simulation of independent paths across multiple processors, achieving near-linear speedup. The FFT-based approach reformulates the convolution operation in Equation (4.3) as:

### v2

@@ 3.2 Proof of main theorem
We proceed with the proof following the aforementioned four steps. First, we need to demonstrate how the convergence target $\Lambda_{t}$ can be expressed in terms of the operator $\mathcal{G}^{\alpha-1}$ defined in (3.3).
@@
The second step, we prove that $Y^{(n)}\Rightarrow Y$ in Hölder space $\mathcal{C}^{\lambda}$ for $\lambda<1/2$. The Donsker invariance principle for convergence in Hölder spaces was first established by lamperti1962convergence. ravckauskas2004necessary provided necessary and sufficient conditions for the convergence of i.i.d. partial sum processes in Hölder spaces, while results for martingale difference sequences were derived by giraudo2016holderian from a dynamical systems perspective. However, these results concern only the scaling limits of a single sequence and do not directly apply to triangular arrays. Actually, FCLT for martingale difference arrays is well established in the Skorokhod space (hall2014martingale). Consequently, once tightness in the Hölder space is verified, the desired convergence in this space follows immediately.
@@
See A.4
Subsequently, we establish the continuity of the operator $\mathcal{G}^{\alpha-1}$, thereby enabling the application of the continuous mapping theorem. Indeed, this result is encompassed within Proposition 2.2 of horvath2024functional, where a proof of continuity for general GFOs is provided. We directly apply their result to obtain the following proposition.
@@
According to Lemma 3.10 in horvath2024functional, we can readily extend this weak convergence to the continuous function space $\mathcal{C}([0,1])$ and Skorokhod space $\mathcal{D}([0,1])$.
We have established the convergence of $\widetilde{\Lambda}_{t}^{(n)}$. However, our primary process of interest is $\Lambda_{t}^{(n)}$ given by (2.5), with $\widetilde{\Lambda}_{t}^{(n)}$ serving merely as an auxiliary process. Finally, we will demonstrate that the two processes are asymptotically indistinguishable. By virtue of Theorem 3.1 in billingsley2013convergence, it follows that $\Lambda_{t}^{(n)}\Rightarrow\Lambda_{t}$.
@@
See A.5
@@ 4.1 Gamma-GED-EGARCH($\infty$) model
If $\varepsilon_{n}$ has unit variance, $\sqrt{\varphi(\lambda_{n})}$ is in fact the conditional volatility of $y_{n}$. Here, the function $\varphi:\mathbb{R}\to\mathbb{R}^{+}$ is monotonic and differentiable, referred to as the link function. Moreover, if $\varphi$ is the identity mapping (restricted to $\mathbb{R}^{+}$), then the time-varying parameter $\lambda_{n}$ directly characterizes the conditional variance of $y_{n}$. According to buccheri2021continuous, the score in this case can be expressed as
@@
Following the terminology in harvey2017volatility, we refer to this model as the Gamma-GED-EGARCH($\infty$). When $\nu=2$, the GED reduces to standard normal distribution, and this model corresponds to the EGARCH model in nelson1991conditional but without asymmetry. Based on the previous definitions, we define the following sequence scaling process:
$\begin{split}&X_{t}^{(n)}=\frac{1}{\sqrt{\chi n}}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}\varepsilon_{i},\quad\varepsilon_{i}|\mathcal{F}_{i-1}\stackrel{{\scriptstyle d}}{{\sim}}\ \text{GED}(\nu),\\ &\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\frac{\nu}{2}|\varepsilon_{n-i}|^{\nu}-1\right],\end{split}$ (4.1)
where $\phi_{i}\sim K/i^{1+\alpha}$, $1-a_{n}\sim n^{-\alpha}$, $\omega_{n}\sim n^{-1/2}$, $\chi=\mathbb{E}(\varepsilon^{2})=2^{2/\nu}\frac{\Gamma(3/\nu)}{\Gamma(1/\nu)}$.
@@
See B.1
The limit (4.2) is actually a rough volatility model, where $\Lambda_{t}$ characterizes the log-volatility of assets log-price $X_{t}$. Note that $\Lambda_{t}$ follows a rough OU process but is different with that in gatheral2018volatility which characterizes log-volatility with a OU process driven by rough fractional Brownian motion. Because the kernel $(t-s)^{\alpha-1}$ not only emerge in noise but also in drift term, this specification is analogous with rough CIR in jaisson2016rough or rough Heston in el2019characteristic.
@@ 4.2 A extension as rough volatility approximations
Before (4.2) can be considered a “useful” rough volatility model, there are still two important issues we need to deal with. First, the initial value of $(X_{t},\Lambda_{t})$ is zero in (4.2); second, the two Brownian motions are independent, which prevents the model from capturing the leverage effect in the market.
For the first issue, we adopt the strategy in wang2025rough, add a baseline in the dynamic of $\lambda_{t}^{(n)}$, that is
@@ 4.3 Numerical simulation
Based on the convergence result, we will simulate the time series (4.3) to approximate rough volatility model (4.4). Accordingly, the simulated paths can be employed to price a variety of options via Monte Carlo methods. In this part, the parameter choices are aligned with those adopted in prior numerical studies, such as callegaro2021fast, ma2022fast and wang2025rough, these values are:
@@
It can be observed that as $\alpha$ increases, the volatility paths exhibit greater smoothness, and the mean-reverting behavior becomes more evident. In other words, a smaller $\alpha$ not only implies rougher paths but also corresponds to a slower mean-reversion, which is consistent with the observations reported in gatheral2018volatility.
@@
Table 1 reports the option prices computed via Algorithm 1 for different values of the roughness parameter $\alpha$, with standard errors shown in parentheses. We set $M=5\times 10^{5}$, $n=500$, $T=1$, with strike prices $K=80,90,100,110,120$. For the barrier options, the barrier level is set to $B_{u}=110$ for the Up-In call and $B_{d}=90$ for the Down-Out put. It can be seen that as $\alpha$ decreases, European calls and puts prices rise, reflecting the fat-tailed return distributions generated by rough volatility processes. Asian options follow similar patterns but with attenuated sensitivity, as price averaging dampens the impact of path roughness. Lookback options show the most pronounced positive sensitivity to roughness, as their payoffs directly depend on extreme price realizations that become more frequent under rough volatility. Barrier options exhibit the most intricate response. While barrier calls increase with roughness, corresponding barrier puts decrease in value. This occurs because roughness simultaneously increases potential terminal payoffs but also raises the probability of early barrier triggering.
Table 1: Option pricing results under the rough volatility model (4.4) for different $\alpha$
Panel A: $\alpha=0.999$
@@
Panel B: $\alpha=0.62$
Strike European Asian Lookback Barrier
Call Put Call Put Call Put Call Put
80 21.4739 1.4614 20.1284 0.1205 37.2465 2.6685 18.6106 0.0000
(0.0274) (0.0060) (0.0166) (0.0012) (0.0213) (0.0077) (0.0298) (0.0000)
90 14.0136 4.0010 11.1613 1.1535 27.2465 7.3286 13.0776 0.0000
(0.0240) (0.0106) (0.0147) (0.0043) (0.0213) (0.0125) (0.0246) (0.0000)
100 8.4253 8.4128 4.7481 4.7402 17.2465 15.4061 8.2971 0.1731
(0.0196) (0.0156) (0.0106) (0.0091) (0.0213) (0.0152) (0.0197) (0.0013)
110 4.6899 14.6774 1.5125 11.5046 9.4320 25.4061 4.6899 1.0888
(0.0150) (0.0203) (0.0061) (0.0134) (0.0187) (0.0152) (0.0150) (0.0047)
120 2.4363 22.4237 0.3641 20.3562 4.8210 35.4061 2.4363 2.9263
(0.0109) (0.0240) (0.0029) (0.0157) (0.0144) (0.0152) (0.0109) (0.0092)
Panel C: $\alpha=0.501$
Strike European Asian Lookback Barrier
Call Put Call Put Call Put Call Put
80 21.4982 1.4699 20.1332 0.1252 37.2660 2.6897 18.6239 0.0000
(0.0274) (0.0060) (0.0166) (0.0012) (0.0214) (0.0077) (0.0298) (0.0000)
90 14.0469 4.0186 11.1787 1.1707 27.2660 7.3630 13.0986 0.0000
(0.0240) (0.0106) (0.0147) (0.0044) (0.0214) (0.0125) (0.0246) (0.0000)
100 8.4553 8.4269 4.7727 4.7647 17.2660 15.4459 8.3241 0.1699
(0.0196) (0.0157) (0.0106) (0.0092) (0.0214) (0.0152) (0.0197) (0.0013)
110 4.7082 14.6799 1.5238 11.5158 9.4538 25.4459 4.7082 1.0710
(0.0150) (0.0203) (0.0061) (0.0134) (0.0188) (0.0152) (0.0150) (0.0046)
120 2.4448 22.4165 0.3717 20.3636 4.8332 35.4459 2.4448 2.8889
(0.0109) (0.0240) (0.0030) (0.0158) (0.0144) (0.0152) (0.0109) (0.0091)
To further verify the validity of our results, we compute the pricing outcomes when $\alpha=1$. In this case, model (4.4) reduces to the classical lognormal stochastic volatility model proposed by scott1987option:
$\begin{split}&{\rm d}X_{t}=-\frac{1}{2}{\rm e}^{2\Lambda_{t}}{\rm d}t+{\rm e}^{\Lambda_{t}}{\rm d}B_{t},\\ &{\rm d}\Lambda_{t}=\kappa(\theta-\Lambda_{t}){\rm d}t+\sigma{\rm d}W_{t},\\ &\text{Cov}(B_{t},W_{t})=\rho t,\end{split}$
where $\sigma=\sqrt{\rho^{2}+\zeta^{2}}\kappa\mu$. Using the same parameter values and Monte-Carlo settings as in the previous experiments, we report the option prices in Table 2. The numbers in parentheses represent the percentage errors relative to the rough volatility model with $\alpha=0.999$. The errors are mostly within 0.5%, with minor deviations attributable to the diminishing roughness. These results confirm that the option prices computed from our model are consistent with those from the classical specification.
Table 2: Option pricing results under the lognormal stochastic volatility model (case $\alpha=1$)
Strike European Asian Lookback Barrier
Call Put Call Put Call Put Call Put
80 21.3770 1.4003 20.1121 0.1086 36.9692 2.5534 18.4349 0.0000
(-0.01%) (-1.08%) (-0.10%) (-2.85%) (0.02%) (-1.08%) (-0.16%) (0.00%)
90 13.8839 3.9071 11.1075 1.1040 26.9692 7.1480 12.9184 0.0000
(-0.08%) (-0.63%) (-0.20%) (-0.37%) (0.03%) (-0.51%) (-0.14%) (0.00%)
100 8.2834 8.3067 4.6752 4.6716 16.9692 15.1966 8.1519 0.1812
(-0.13%) (-0.30%) (-0.38%) (0.00%) (0.05%) (-0.23%) (-0.16%) (1.27%)
110 4.5550 14.5783 1.4563 11.4527 9.1719 25.1966 4.5550 1.1266
(0.01%) (-0.09%) (-0.34%) (0.11%) (0.21%) (-0.14%) (0.01%) (1.38%)
120 2.3233 22.3466 0.3396 20.3360 4.6129 35.1966 2.3233 3.0172
(0.43%) (-0.02%) (0.80%) (0.10%) (0.55%) (-0.10%) (0.43%) (1.06%)
We then examine the shape of the implied volatility surface generated by the model, focusing specifically on the at-the-money (ATM) volatility skew $\psi(\tau)$, defined as
$\psi(\tau):=\left|\frac{\partial\sigma_{IV}(k,\tau)}{\partial k}\right|_{k=0},$
where $k=\log(S_{0}/K)$ denotes the log-moneyness and $\tau$ represents the time to maturity.
Empirical evidence suggests that the volatility skew exhibits a power-law decay as time to maturity increases (gatheral2018volatility). Theoretical work by fukasawa2011asymptotic established that when volatility is driven by fractional Brownian motion with Hurst parameter $H$, the ATM skew asymptotically follows $\psi(\tau)\sim\tau^{H-1/2}$ for small $\tau$. We compute the ATM volatility skew generated from our model, as shown in Figure 2. To assess the power-law relationship, we employ a double logarithmic scale and perform a linear regression on the log-transformed data. The results largely validate a power-law behavior ($R^{2}>0.85$), with two observations consistent with prior studies: 1. Rougher volatility paths generate steeper initial skews that subsequently decay more rapidly as $\tau$ increases. This aligns with the notion that greater path roughness implies a stronger local impact of volatility shocks, resulting in a more pronounced skew effect, especially for short maturities; 2. In the limit case where roughness is nearly absent ($\alpha=0.999$), the volatility skew is observed to approximate a constant across maturities.
Figure 2: Power-law relationship between ATM volatility skew and time to maturity (Log-Log plot)
Finally, we are concerned with the computational speed of this approximation. To accelerate the path generation process, we employ two techniques: parallel computing and FFT-based convolution. Parallel computing distributes the simulation of independent paths across multiple processors, achieving near-linear speedup. The FFT-based approach reformulates the convolution operation in Equation (4.3) as: